# Segre products of cluster algebras

### v1

@@ 2.2 Relationship with the Segre product
$d=\sum\limits_{B^{1}_{jk}>0}B^{1}_{jk}G_{j}^{1}=\sum\limits_{B^{1}_{jk}<0}-B^{1}_{jk}G_{j}^{1}$
noting that the two are equal since $B_{1}^{T}G_{1}=0$ and $\deg\mu_{k}(x_{k})=d-\deg x_{k}$.
An analogous argument shows that $\varphi(\mu_{k}(y_{k}))=x_{s_{1}}^{\deg\mu_{k}(y_{k})}\otimes\mu_{k}(y_{k})$ for $y_{k}\in\underline{y}$.
Now, let $\underline{z}=(z_{1},\dotsc,z_{n_{1}+n_{2}-1})$ be a cluster $m$ mutation steps away from the initial cluster $\underline{x}$, i.e. $\underline{z}=\mu_{\underline{p}}(\underline{x})$ for some mutation path $\underline{p}$ of length $m$, and assume that
$\varphi(z_{j})=z_{j}\otimes y_{s_{2}}^{\deg z_{j}}$
for $z_{j}\in\mathcal{A}_{1}$. Denote by $B^{\prime}$ the corresponding exchange matrix. We claim that, for $z_{k}\in\mathcal{A}_{1}$,
$\varphi(\mu_{k}(z_{k}))=\mu_{k}(z_{k})\otimes y_{s_{2}}^{\deg\mu_{k}(z_{k})}.$
Indeed, the same calculation as above with $\underline{z}$ and $B^{\prime}$ shows that this is the case and similarly $\varphi(\mu_{k}(z_{k}))=x_{s_{1}}^{\deg\mu_{k}(z_{k})}\otimes\mu_{k}(z_{k})$, when $z_{k}\in\mathcal{A}_{2}$.
The above allows us to define $\varphi\colon\mathcal{A}_{1}\Box\mathcal{A}_{2}\to\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ on the generating set of cluster variables as above. We have seen that this respects the defining (exchange) relations and can therefore be extended to an algebra homomorphism. It is clearly injective on the generating set of the domain and therefore injective. ∎
Note that
$\varphi(x_{j})$ $=x_{j}\otimes y_{s_{2}}^{\deg x_{j}}\in(\mathcal{A}_{1})_{\deg x_{j}}\otimes(\mathcal{A}_{2})_{\deg x_{j}\deg y_{s_{2}}}$
$\varphi(y_{j})$ $=x_{s_{1}}^{\deg y_{j}}\otimes y_{j}\in(\mathcal{A}_{1})_{\deg y_{j}\deg x_{s_{1}}}\otimes(\mathcal{A}_{2})_{\deg y_{j}}$
$\varphi(x_{s})$ $=x_{s_{1}}\otimes y_{s_{2}}\in(\mathcal{A}_{1})_{\deg x_{s_{1}}}\otimes(\mathcal{A}_{2})_{\deg y_{s_{2}}}$
and so we do not land in the Segre product without extra constraints. However, we also see that the obstacle is the degree of the gluing frozen variables: fixing these to be 1, we immediately obtain that the map is a graded homomorphism to the Segre product.
Proposition 2.7.
Let $\mathcal{A}_{i}=(\underline{x}_{i},\underline{\textup{ex}}_{i},B_{i},G_{i})$, $i=1,2$ be graded cluster algebras such that there exist $s_{1}\in\{1,\dotsc,n_{1}\}\setminus\underline{\textup{ex}}_{1}$ and $s_{2}\in\{1,\dotsc,n_{2}\}\setminus\underline{\textup{ex}}_{2}$ of degree 1.
$\varphi(x_{j})$ $=x_{j}\otimes y_{s_{2}}^{\deg x_{j}}\quad\text{for }j\in\{1,\dotsc,n_{1}\},$
$\varphi(y_{j})$ $=x_{s_{1}}^{\deg y_{j}}\otimes y_{j}\quad\text{for }j\in\{1,\dotsc,n_{2}\}\ \text{and}$
$\varphi(x_{s})$ $=x_{s_{1}}\otimes y_{s_{2}}$
is an injective graded algebra homomorphism, with the property that the above formulæ hold for any cluster of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$. ∎
It remains to show that $\varphi$ is an isomorphism, for then the cluster structure defined on $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ indeed induces a cluster structure on the Segre product $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$. This again requires a further assumption.
Definition 2.8.
Let $\mathcal{A}=(\underline{x},\underline{\textup{ex}},B,G)$ be a graded cluster algebra. We say that a cluster $\underline{y}$ of $\mathcal{A}$ is homogeneous of degree $d$ if all cluster variables in $\underline{y}$ have the same degree $d$.
Theorem 2.9.
Let $\mathcal{A}_{i}=(\underline{x}_{i},\underline{\textup{ex}}_{i},B_{i},G_{i}=\mathds{1})$, $i=1,2$ be graded cluster algebras such that $\underline{x}_{1}$ and $\underline{x}_{2}$ are homogeneous of degree 1 and fix $s_{1}\in\{1,\dotsc,n_{1}\}\setminus\underline{\textup{ex}}_{1}$ and $s_{2}\in\{1,\dotsc,n_{2}\}\setminus\underline{\textup{ex}}_{2}$.
$\varphi(x_{j})$ $=x_{j}\otimes y_{s_{2}}\quad\text{for }j\in\{1,\dotsc,n_{1}\},$
$\varphi(y_{j})$ $=x_{s_{1}}\otimes y_{j}\quad\text{for }j\in\{1,\dotsc,n_{2}\}\ \text{and}$
$\varphi(x_{s})$ $=x_{s_{1}}\otimes y_{s_{2}}$
Thus the construction above endows $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$ with the structure of a cluster algebra.
Proof.
@@
$z_{1}\otimes z_{2}$ $=(z_{1}\otimes y_{s_{2}})(x_{s_{1}}\otimes z_{2})(x_{s_{1}}^{-d}\otimes y_{s_{2}}^{-d})=\varphi(z_{1})\varphi(z_{2})\varphi(x_{s})^{-d}.$
Hence, $\operatorname{Im}\varphi$ contains a generating set for $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$, and so $\varphi$ is surjective. The claim follows. ∎
@@
Notice that in proving surjectivity, we required $\varphi(x_{s})=x_{s_{1}}\otimes y_{s_{2}}$, and hence $x_{s_{1}}$ and $y_{s_{2}}$ themselves, to be invertible, but no other frozen variables needed to be invertible for the proof to hold.
The following is now immediate from Lemmas 2.3 and 2.4.
Corollary 2.11.
Let $\mathcal{A}_{i}=(\underline{x}_{i},\underline{\textup{ex}}_{i},B_{i},G_{i}=\mathds{1})$, $i=1,2$ be graded cluster algebras such that $\underline{x}_{1}$ and $\underline{x}_{2}$ are homogeneous of degree 1 and fix $s_{1}\in\{1,\dotsc,n_{1}\}\setminus\underline{\textup{ex}}_{1}$ and $s_{2}\in\{1,\dotsc,n_{2}\}\setminus\underline{\textup{ex}}_{2}$.
Then
  1. (i)
    the cluster algebras $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$ and $\mathcal{A}_{2}\overline{\otimes}\mathcal{A}_{1}$ are isomorphic as cluster algebras;
  2. (ii)
    $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$ is of finite type if and only if $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are;
  3. (iii)
    writing $\kappa(\mathcal{A})$ for the number of cluster variables of a cluster algebra $\mathcal{A}$, we have $\kappa(\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2})=\kappa(\mathcal{A}_{1})+\kappa(\mathcal{A}_{2})-1$ when these numbers are all finite; and
  4. (iv)
    writing $K(\mathcal{A})$ for the number of clusters of $\mathcal{A}$, we have $K(\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2})=K(\mathcal{A}_{1})K(\mathcal{A}_{2})$ when these numbers are all finite. ∎
@@
In the above theorem, we require the input clusters to be homogeneous of degree one. This was necessary to ensure that the image of $\varphi$ generates the Segre product as defined. We note that this condition is very restrictive and it would be desirable for it to be weakened. However, without it, it does not seem feasible to describe the image in as simple a fashion as for the standard Segre product.
We also observe that, as a result, the construction of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ and hence the cluster structure induced on $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$ is strongly “rooted”, i.e. dependent on a choice of initial clusters with specific properties. Even in a graded cluster algebra that has a homogeneous cluster, only with very specific exchange matrices we will find that other clusters are also homogeneous.

### v2

@@ 2.2 Relationship with the Segre product
$d$ $=\sum_{x_{j}}[C_{x_{j},x_{k}}]_{+}\deg x_{j}=\sum_{C_{x_{j},x_{k}}>0}C_{x_{j},x_{k}}H_{x_{j}}$
$=\sum_{C_{x_{j},x_{k}}<0}-C_{x_{j},x_{k}}H_{x_{j}}=\sum_{x_{j}}[C_{x_{j},x_{k}}]_{-}\deg x_{j}$
noting that the third equality holds since $C^{T}H=0$. Also, we use that $\deg\mu_{k}(x_{k})=d-\deg x_{k}$.
Note that the fifth equality is where the assumption that $\deg x=1$ is used: without it, the claimed equality of $d$ with the other stated quantities need not hold.
An analogous argument shows that $\varphi(\mu_{k}(y_{k}))=x^{\deg\mu_{k}(y_{k})}\otimes\mu_{k}(y_{k})$ for $y_{k}\in\underline{y}_{2}$ mutable, noting that this time, it is $\deg y=1$ that is required.
Since we have $\deg x=\deg y=1$, the above tells us that for any cluster variable $x^{\prime}$ of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$, we either have $\varphi(x^{\prime})=x^{\prime}\otimes y^{\deg x^{\prime}}$ or $\varphi(x^{\prime})=x^{\deg x^{\prime}}\otimes x^{\prime}$ and hence $\varphi(x^{\prime})\in(\mathcal{A}_{1})_{\deg x^{\prime}}\otimes(\mathcal{A}_{2})_{\deg x^{\prime}}$. That is, the image of $\varphi$ is contained in the Segre product $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$ without any further constraints and the map $\varphi$ is a graded map.
@@
$z_{1}\otimes z_{2}$ $=(z_{1}\otimes y^{d})(x^{d}\otimes z_{2})(x^{-d}\otimes y^{-d})=\varphi(z_{1})\varphi(z_{2})\varphi(z)^{-d}.$
Hence, $\operatorname{Im}\varphi$ contains a generating set for $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$, and so $\varphi$ is surjective onto $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$. The claim follows. ∎
@@
One might be tempted to try changing the specification of the map $\varphi$ to
$\varphi(x_{j})$ $=x_{j}^{\deg y}\otimes y^{\deg x_{j}}\quad\text{for }x_{j}\in\underline{x}_{1}\setminus\{x\},$
$\varphi(y_{j})$ $=x^{\deg y_{j}}\otimes y_{j}^{\deg x}\quad\text{for }y_{j}\in\underline{x}_{2}\setminus\{y\}\ \text{and}$
$\varphi(z)$ $=x^{\deg y}\otimes y^{\deg x}$
in an attempt to avoid the $\deg x=\deg y=1$ assumption. Note that one should however ask for $\deg x$ and $\deg y$ strictly positive, to avoid issues with needing inverses of arbitrary cluster variables.
While this does indeed fix the issue with $d$ that occurs in the calculation in the above proof for $x_{k}\in\underline{y}_{1}$, the appearance of $x^{\deg y}$ in the first tensor factor means that we do not obtain $\mu_{k}(x_{k})$ unless $\deg y=1$.
More explicitly, following the same approach as in the previous proof, one would arrive at
$\frac{1}{x_{k}^{\deg y}\otimes y^{\deg x_{k}}}\Bigg{[}\prod\limits_{x_{j}\in\underline{y}_{1}}x_{j}^{[C_{x_{j},x_{k}}]_{+}\deg y}\otimes y^{d}+\prod\limits_{x_{j}\in\underline{y}_{1}}x_{j}^{[C_{x_{j},x_{k}}]_{-}\deg y}\otimes y^{d}\Bigg{]}$
but this is not equal to
$\frac{1}{x_{k}^{\deg y}}\Bigg{(}\prod\limits_{x_{j}\in\underline{y}_{1}}x_{j}^{[C_{x_{j},x_{k}}]_{+}}+\prod\limits_{x_{j}\in\underline{y}_{1}}x_{j}^{[C_{x_{j},x_{k}}]_{-}}\Bigg{)}^{\deg y}\otimes y^{d-\deg x_{k}}$
if $\deg y\neq 1$.
By symmetry, the other case tells us that we also need $\deg x=1$. That is, the degree 1 assumption is unavoidable.
@@
Notice that in proving surjectivity, we required $\varphi(z)=x\otimes y$, and hence $x$ and $y$ themselves, to be invertible, but no other frozen variables needed to be invertible for the proof to hold.
Remark 2.10.
Via Lemma 2.4, we see that the cluster structure on $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ and hence that on $\mathcal{A}_{1}\overline{\otimes}\mathcal{A}_{2}$ is independent of the choices of initial seeds. Therefore the only requirements to obtain a cluster structure on the Segre product are the existence of a frozen variable of degree 1 for each factor.
Graded cluster algebras with at least one frozen variable of degree one are, perhaps surprisingly, ubiquitous. Many examples arising geometrically have this property: coordinate rings of Grassmannians and more generally partial flag varieties and their cells ([GLS11]), double Bruhat cells ([BFZ05]) and, as motivated this work, positroid varieties ([GL19]).
Note too that the claims on the cluster structure of $\mathcal{A}_{1}\Box\mathcal{A}_{2}$ in Corollary 2.5 therefore also apply to the induced cluster structure on the Segre product.